OpenStudy (anonymous):

4x-7/2 is greater than 4@petiteme

OpenStudy (petiteme):

You there?

OpenStudy (anonymous):

yes

OpenStudy (petiteme):

First, transpose -7/2 to the other side.

OpenStudy (petiteme):

It will be 4x > 4 + 7/2 right?

OpenStudy (anonymous):

ok

OpenStudy (petiteme):

Then both should be fraction..So that you can add them. Can you do it?

OpenStudy (anonymous):

not clear

OpenStudy (petiteme):

It will be like this...

OpenStudy (petiteme):

\[4x > \frac{ 6 }{ 2 } + \frac{ 7 }{ 2 }\]

OpenStudy (petiteme):

Do you know how I got it? :D

OpenStudy (anonymous):

not really i was thing 4=8/2

OpenStudy (petiteme):

Huh?

OpenStudy (petiteme):

So 4x > 13/2 right?

OpenStudy (anonymous):

but how did you get the 6/2

OpenStudy (petiteme):

The GCF is 2 right?

OpenStudy (petiteme):

So it is 4/1 .... 2 is multiplied to 1 then added to 4

OpenStudy (petiteme):

You will get 6/2

OpenStudy (anonymous):

i see

OpenStudy (petiteme):

Then I can know add it cause they both have the same denominator. So it is 13/2 right?

OpenStudy (petiteme):

now*

OpenStudy (petiteme):

So the equation is now 4x > 13/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (petiteme):

Divide both sides by 4

OpenStudy (anonymous):

x=13/8 or 1.625

OpenStudy (petiteme):

I think you did it wrong. You should multiply 4 to 13

OpenStudy (petiteme):

You should get 52/2

OpenStudy (anonymous):

how

OpenStudy (anonymous):

52/2 is 4.13, not 13/2 divided by 4

OpenStudy (petiteme):

Ohhh yeah my bad :/ Hehe. So it will be 13/8 ^_^

OpenStudy (petiteme):

You got it right ^_^ Good job!

OpenStudy (anonymous):

ok

OpenStudy (ikram002p):

4x-7/2 > 4 4x-7>8 4x>13 x>13/4

OpenStudy (petiteme):

Close this question if you don't have any questions anymore:D

OpenStudy (anonymous):

can I ask one more please!

OpenStudy (petiteme):

Sure...

OpenStudy (anonymous):

by the way how do u write the symbols like greater than

OpenStudy (petiteme):

In your keyboard :) Click shift then period

OpenStudy (anonymous):

-6\[-6<-2x+6\le12\]

OpenStudy (petiteme):

Now subtract -2 from each part

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[-8<-2x +4\le10\]

OpenStudy (petiteme):

NO WAIT. Subtract 6 to both parts-.- Sorry.

OpenStudy (petiteme):

Not -2 :/ Sorry.

OpenStudy (anonymous):

ok

OpenStudy (petiteme):

It is -12 < -2x < 6

OpenStudy (anonymous):

\[that would be -12<-2x \le6\]

OpenStudy (petiteme):

Correct! Then divide both sides by -2 :)

OpenStudy (petiteme):

Because you are dividing a negative number, the inequalities change direction.

OpenStudy (anonymous):

\[-6>x \ge-3\]

OpenStudy (petiteme):

The -6 should only be 6 :)

OpenStudy (petiteme):

Negative divided by a negative is a positive ^_^

OpenStudy (anonymous):

oh yea..thanks

OpenStudy (petiteme):

YW ^_^

OpenStudy (anonymous):

\[6>x \ge-3 \right?\]

OpenStudy (anonymous):

\[6>x \ge-3\]

OpenStudy (anonymous):

right?

OpenStudy (petiteme):

Good Job! :)

OpenStudy (anonymous):

is that the final answer?

OpenStudy (petiteme):

But to be neat it is better to have the smaller number on the left, larger on the right. So let us swap them over and make sure the inequalities point correctly.

OpenStudy (anonymous):

ok

OpenStudy (petiteme):

:)

OpenStudy (anonymous):

\[-3lex <6\]

OpenStudy (anonymous):

\[-3\le x <6\]

OpenStudy (anonymous):

is that right?

OpenStudy (petiteme):

yes :)

OpenStudy (anonymous):

ok

OpenStudy (wolfe8):

Petite is very sweet girl, and helpful too. :)